OpenStudy (anonymous):

The total snowfall in Buffalo, NY one year was 120 inches. If the total snowfall the following year was 90 inches, what was the percent change? Show and explain your solution

OpenStudy (anonymous):

120=100% 120/100=90/x x=90*100/120=75%

OpenStudy (amistre64):

find the difference in the snow fall amounts, and divide it by the original amount

OpenStudy (amistre64):

(120-90)/120

OpenStudy (amistre64):

the first poster seems to be demonstrating that 90 is 75% of 120; the change therefore is the difference between 100% and 75%

OpenStudy (anonymous):

25?

OpenStudy (amistre64):

correct

OpenStudy (anonymous):

how do i explain that?

OpenStudy (amistre64):

prolly by wording how i explained it in your own terms

OpenStudy (anonymous):

k thanks

OpenStudy (amistre64):

90 is 75% or 120; the change from 100% to 75% is 25%

OpenStudy (amistre64):

*75% of 120

OpenStudy (anonymous):

thnaks

OpenStudy (anonymous):

*thanks